OpenStudy (anonymous):

(2/5x)^(5/6) Simplify

OpenStudy (anonymous):

\[\left(\begin{matrix}2 \\ 5\end{matrix}\right)^{5/6}\left(\begin{matrix}1 \\ x ^{}\end{matrix}\right)^{5/6}\]